40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 7 14:04:48 2011, steve, Update, SUS, He/Ne intensity noise effect on oplevs 7x
    Reply  Fri Oct 7 23:23:05 2011, kiwamu, Update, SUS, SRM oplev was oscillating 
       Reply  Wed Oct 19 10:49:35 2011, kiwamu, Update, SUS, SRM oscillation removed SRMoplevKWMtune.png
    Reply  Wed Oct 12 08:43:30 2011, steve, Update, SUS, ITMX oplev improved a bit ITMXoplev.pngITMXoplevservo_ON.png
       Reply  Thu Oct 13 14:23:09 2011, steve, Update, SUS, ITMX oplev with 3 mm beam on qpd ITMXoplev1103p#2.pngITMXoplev.png
          Reply  Mon Oct 17 17:07:30 2011, steve, Update, SUS, ETMX oplev returning beam od 3 mm ETMX-oplev.jpg
             Reply  Wed Oct 19 09:44:10 2011, steve, Update, SUS, ETMX oplev power spectrum ETMXoplev2lens.png
             Reply  Fri May 18 15:25:15 2012, steve, Update, SUS, AR coated lens swapped in at ETMX oplev ETMXoplevlensAR.png
Message ID: 5630     Entry time: Fri Oct 7 14:04:48 2011     Reply to this: 5636   5655
Author: steve 
Type: Update 
Category: SUS 
Subject: He/Ne intensity noise effect on oplevs 

The SRM bounce peak 18.33 Hz. Suresh helped me to retune filter through Foton, but we failed to remove it.

ITMX_OPLEV_PERROR shows high coherence with oplev laser. This is our lousiest set up. I will work on it next week.

Generally speaking we can say that JDSU-Uniphase 1103P and 1125P laser intensity noise do not limit oplev servo performance.

However, the overall well being of filters, gain settings, beam sizes on qpds are poor.


Attachment 1: PRMoplevINTn.png  51 kB  | Hide | Hide all
Attachment 2: BSoplevINTn.png  51 kB  | Hide | Hide all
Attachment 3: ITMXoplevINTn.png  52 kB  | Hide | Hide all
Attachment 4: ITMYoplevINTn.png  53 kB  | Hide | Hide all
Attachment 5: SRM_oplevINTn.png  53 kB  | Hide | Hide all
Attachment 6: ETMYoplevINTn3.png  53 kB  | Hide | Hide all
Attachment 7: ETMXoplevINTn.png  52 kB  | Hide | Hide all
ELOG V3.1.3-